Solve 2sin x - cos^2 x = 2 ober the domain 0 < x < 2pie
Answer by stanbon(75887)

Solve
2sin x - cos^2 x = 2 over the domain 0 < x < 2pie
-----
2sin(x) - (1-sin^2(x)) = 2
---
sin^2(x) + 2sin(x) -3 = 0
------
Factor:
(sin(x)+3)(sin(x)-1) = 0
---
sin(x) = -3 or sin(x)= 1
---
x is extraneous in sin(x) = -3
----------------------------------
x = pi/2 in sin(x) = 1
=========================
Cheers,
Stan H.